Given a normal distribution with a mean of 200 and the deviation is10. Find the area above 188.



1
Expert's answer
2022-04-12T16:28:43-0400

"P(X>188)=P(Z>\frac{188-200}{10})=P(Z>-1.2)=1-P(Z<-1.2)=0.8849."